### v1

# Computational techniques for sheaf cohomology of locally profinite sets

Abstract.
We compute the sheaf cohomology with constant $\mathbb{Z}_{2}$ coefficients of a concrete class of locally profinite sets of independent interest. We introduce $k$-Fubini partitions to aid in constructions, which witness a failure of a Fubini theorem analog for these spaces. It is also shown that questions of intermediate cohomology degrees can be reduced to questions about top cohomology degrees by exhibiting nontrivial top cocycles as pointwise limits of coboundaries.
1. Introduction
In this paper we describe a toolbox for computing sheaf cohomology groups of certain “infinitary cuboid” spaces. These are locally profinite sets, in that they arise from the deletion of a point from a profinite set. Profinite sets [^1] More concisely, Stone spaces; more verbosely, totally disconnected compact Hausdorff spaces., in turn, are important for their role in Stone duality, where they form one end of a logical-topological correspondence with Boolean rings.
Recent work has demonstrated that analysis of the cohomology of locally profinite sets, ported across Stone duality, yields novel information about the corresponding rings. Specifically, the construction of a locally profinite set of size $\aleph_{n}$ and cohomological dimension $n$, due to Aoki [1], yields a nondescendable faithfully flat map of Boolean rings of size $\aleph_{\omega}$. This, in turn, has implications for the theory of light condensed sets, due to Clausen and Scholze [2, 3]. Aoki’s constructions are infinitary-combinatorial in nature; the essential goal of this paper is to further those arguments by providing novel tools and techniques for computing the cohomology of locally profinite sets. Nevertheless, our methods remain largely elementary, and no forcing machinery is involved.
The paper contains three main parts, corresponding to the three main tools developed. After expositing the spaces under study in section 2, in section 3 we analyze their sheaf cohomology directly by exhibiting nontrivial cocycles via a nonconstructive argument. Here our arguments are within ZFC, but our conclusions are strengthened by cardinal arithmetic bounds. In section 4, we define $k$-Fubini partitions of the spaces, which witness a failure of a certain analog of Fubini’s theorem. We use these partitions to show both vanishing and nonvanishing of certain cohomology groups; we also prove a converse to one of Aoki’s main results. Lastly, in section 5 we characterize nonvanishing cohomology purely in terms of nonvanishing top cohomology. A better understanding of these top cohomology groups is thus sufficient to resolve all remaining cases. All our arguments may be construed as taking place in ZF, with choice only needed in a few places, such as in the existence of an injection $\aleph_{1}\hookrightarrow 2^{\aleph_{0}}$.
I am grateful to Justin Tatch Moore and Jeffrey Bergfalk for invaluable advising, as well as to the NSF Graduate Research Fellowship for funding the period of my graduate study during which the majority of this research was undertaken.
2. Definitions and overview of results
Much of our notation is standard:
  • Tuples of cardinals, ordinals, etc. are denoted by $\vec{\kappa},\vec{\lambda},\vec{\beta},\vec{x}$, etc. (We reserve $\alpha$ to refer to the one-point compactification $\alpha(X)$ of a space $X$.)
  • We frequently denote by $n+1$ the set $\{0,\ldots,n\}$, and given $j\in n+1$ and a tuple $\vec{x}$ we write $\vec{x}^{j}$ for the tuple of length $n$ obtained by deleting the $j$-th coordinate from $\vec{x}$. We also extend this notation as follows: for $A\subseteq n+1$, the expression $\vec{x}^{A}$ refers to the tuple obtained by deleting all elements whose index lies in $A$. By analogy, $\vec{x}\upharpoonright_{A}$ refers to the tuple of elements of $\vec{x}$ whose indices lie in $A$.
  • The expressions $A]^{k},[A]^{\geq k}$ refer to the sets of subsets of $A$ of size $k$ and of size at least $k$, respectively.
  • For tuples $\vec{\lambda},\vec{\kappa}$, the relation $\vec{\lambda}\leq\vec{\kappa}$ will always refer to the coordinate-wise partial order.
Definition 2.1.
Let $\vec{\kappa}$ be a weakly increasing tuple of cardinals. The space $X(\vec{\kappa})$ is defined
$X(\vec{\kappa}):=\prod_{i\leq n}\alpha(\kappa_{i}),$
where as above $\alpha(\kappa_{i})=\kappa_{i}\amalg\{\infty\}$ denotes the one-point compactification of the discrete space of size $\kappa_{i}$. We denote by $X(\vec{\kappa})^{-}$ the subspace obtained by deleting the point $(\infty,\ldots,\infty)$.
Study of the $X(\vec{\kappa})^{-}$ prior to Aoki largely took place in the 1960s; the space $X(\aleph_{0},\mathfrak{c})^{-}$ is an important counterexample in point-set topology, where it’s known as Thomas’ plank [4]. Wiegand [5, 6] also considered the spaces $X(\aleph_{0},\aleph_{1},\ldots,\aleph_{n})^{-}$, which Aoki showed have nonzero cohomology in the $n$-th degree.
Remark 2.2.
Our notation differs slightly from Aoki’s; for instance, where he writes $X(0,1,2)$ we write $X(\aleph_{0},\aleph_{1},\aleph_{2})$. We will still typically consider the case where $\sup\vec{\kappa}<\aleph_{\omega}$.
A Leray cover for $X(\vec{\kappa})^{-}$, i.e., one in which all finite intersections are acyclic, is given by the sets $\{\vec{x}:x_{i}\neq\infty\}$ for each $i\leq n$. Aoki used this cover to characterize $H^{\bullet}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ via the associated Čech complex. We introduce it via the following notation:
Definition 2.3.
Let $A\subseteq n+1$, and define
$\mathcal{D}_{A}(\vec{\kappa})$ $=\{\langle x_{i}\rangle\in X(\vec{\kappa}):\forall i\in A,x_{i}\neq\infty\},$
$\mathcal{C}_{A}(\vec{\kappa})$ $=\{f:\mathcal{D}_{A}(\vec{\kappa})\to\mathbb{Z}_{2}\text{ continuous}\},$
where
  • $\mathcal{D}_{A}{(\vec{\kappa})}$ is given the subspace topology;
  • $\mathbb{Z}_{2}$ is given the discrete topology; and
  • $\mathcal{C}_{A}{(\vec{\kappa})}$ inherits the ring structure from $\mathbb{Z}_{2}$, and is given the compact-open topology.
We omit the tuple $\vec{\kappa}$ from the notation when clear, writing $\mathcal{D}_{A},\mathcal{C}_{A}$.
Proposition 2.4 ([1]).
The sheaf cohomology of $X(\vec{\kappa})^{-}$ with constant $\mathbb{Z}_{2}$ coefficients is computed by the Čech complex whose $k$-th term is
$\check{C}^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=\prod_{A\in[n+1]^{k+1}}\mathcal{C}_{A}(\vec{\kappa}).$
More concretely:
  • a $k$-cochain is (represented by) a tuple of functions $\vec{f}=\langle f_{A}:A\in[n+1]^{k+1}\rangle,$ where each $f_{A}:\mathcal{D}_{A}(\vec{\kappa})\to\mathbb{Z}_{2}$ is continuous; i.e., $f_{A}$ is an ordinary function whose support is a clopen subset of $\mathcal{D}_{A}$.
  • A tuple $\vec{f}$ represents a $k$-cocycle if, for any $A^{\prime}\in[n+1]^{k+2}$, we have
    $(d\vec{f})_{A^{\prime}}:=\sum_{i\in A^{\prime}}f_{A^{\prime}\setminus\{i\}}=0.$
  • A $k$-coboundary arises from some tuple $\langle g_{A^{\prime}}:A^{\prime}\in[n+1]^{k}\rangle$ via
    $f_{A}=\sum_{i\in A}g_{A\setminus\{i\}}.$
Aoki notes that Proposition 2.4 immediately implies that the cohomological dimension of $X(\vec{\kappa})^{-}$ is at most $n+1$; moreover, he uses this characterization to show that $H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$ whenever $\kappa_{i}\geq\aleph_{i}$ for all $i\leq n$. Here, a top cocycle is just a function $f:\prod_{i\leq n}\kappa_{i}\to\mathbb{Z}_{2}$, and $f$ represents a coboundary if and only if it is the sum of maps $f_{i}:\prod_{i\leq n}\kappa_{i}\to\mathbb{Z}_{2},i\leq n$, such that $f_{i}$ is mod-finite independent of the $i$-th coordinate.
Our first main tool, proved using this characterization, is the following.
Theorem A.
Let $0<k<n\in\mathbb{N}$, and suppose $\vec{\kappa}=\langle\kappa_{0}\leq\kappa_{1}\leq\cdots\leq\kappa_{n}\rangle$ is a weakly increasing tuple of cardinals such that
$\kappa_{n-k}<\operatorname{cf}\kappa_{n}\quad\text{ and }\quad 2^{\kappa_{0}}\geq\kappa_{n-k+1},$
then $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$.
As a simple corollary, we obtain ZFC “reference spaces” for nonvanishing $H^{k}$ among the $X(\vec{\kappa})^{-}$, namely those indexed by tuples of the form
$\vec{\kappa}=\langle\underbrace{\aleph_{0},\aleph_{0},\ldots,\aleph_{0}}_{n-k+1},\underbrace{\aleph_{1},\aleph_{1},\ldots,\aleph_{1}}_{k}\rangle.$
In section 4, we describe techniques for eliminating the cardinal arithmetic assumptions of Theorem A in certain situations. The main tool is a partition of the product $\prod_{i\leq n}\kappa_{i}$ into $n+1$ pieces, each of which is “thin” in one of the coordinate directions. The situation resembles a failure of Fubini’s theorem, so we call such partitions of the spaces “$k$-Fubini”.
We use this machinery to obtain both vanishing results and embeddings of cohomology groups for certain tuples:
Theorem B.
Let $\vec{\kappa}$ be a length-$(n+1)$ tuple of cardinals which admits a $k$-Fubini partition. Then:
  1. (1)
    $H^{k+1}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0,$ and
  2. (2)
    For any $\vec{\lambda}\leq\vec{\kappa}$ coordinatewise, there is an embedding H^k(X(λ)^-,Z_2) H^k(X(κ)^-,Z_2)
We determine exactly when $k$-Fubini partitions exist, in the process deriving a converse to one of Aoki’s results:
Corollary 2.5.
Let $\vec{\kappa}$ be a tuple of cardinals such that $\kappa_{i}<\aleph_{i}$ for some $i\leq n$. Then $H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$.
Lastly, in the final section we prove the following sufficient condition for nonvanishing cohomology.
Theorem C.
Suppose there exists $B\in[n+1]^{k+1}$ and a continuous map $\varphi:X(\vec{\kappa}^{B})\to H^{k}(X(\vec{\kappa}\upharpoonright B)^{-},\underline{\mathbb{Z}_{2}})$ such that
  1. (1)
    $\varphi\upharpoonright X(\vec{\kappa})^{-}$ is identically zero, and
  2. (2)
    $\varphi(\infty,\ldots,\infty)\neq 0$.
Then $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$.
Theorem C can be read in the following manner: let $\vec{f}=\varphi(\infty,\ldots,\infty)$. Then $\vec{f}$ is nonzero in $H^{k}(\vec{\kappa}\upharpoonright_{B})$, but the map $\varphi$ exhibits $\vec{f}$ as a limit of coboundaries. Theorem C makes use of this incompactness property to construct a nonzero cocycle in the intermediate cohomology of the larger space $X(\vec{\kappa})^{-}$.
3. Proof of Theorem A
We begin with an analysis of the cardinality and relationships between the $\mathcal{C}_{A}(\vec{\kappa})$, for varying $A$ and $\vec{\kappa}$. The following two facts will be useful; their proofs are routine.
Facts.
  1. (1)
    If $A\subseteq A^{\prime}$, then $\mathcal{D}_{A^{\prime}}\subseteq\mathcal{D}_{A}$ is open dense, so that the map $f\mapsto f\restriction\mathcal{D}_{A}$ defines a dense embedding of topological rings $\mathcal{C}_{A}\hookrightarrow\mathcal{C}_{A^{\prime}}$. In light of this and related observations, we will frequently abuse notation when dealing with functions defined on different domains; for instance, given functions $f_{A}\in\mathcal{D}_{A}$ and $f_{B}\in\mathcal{D}_{B}$, we write $f_{A}+f_{B}$ for the function defined on the intersection $\mathcal{D}_{A}\cap\mathcal{D}_{B}=\mathcal{D}_{A\cup B}$.
  2. (2)
    If $j\in A$, then a continuous map $f\in\mathcal{C}_{A}(\vec{\kappa})$ corresponds uniquely to a $\kappa_{j}$-tuple of maps $f_{\beta}\in\mathcal{C}_{A\setminus\{j\}}(\vec{\kappa}^{j})$, and this correspondence in fact defines an isomorphism of topological rings C_A(→κ) ≅(C_A ∖{j}(→κ^j))^κ_j More generally, any inclusion $A\subseteq A^{\prime}$ yields an isomorphism of topological rings C_A’(→κ) ≅C_A’ ∖A(→κ^A)^∏_j ∈Aκ_j.
As a first application of these observations, we find the cardinality of $\mathcal{C}_{A}(\vec{\kappa})$ for arbitrary $A,\vec{\kappa}$.
Proposition 3.1.
Suppose $\vec{\kappa}=\langle\kappa_{0}\leq\ldots\leq\kappa_{n}\rangle$ is a weakly increasing $(n+1)$-tuple of cardinals, and let $A\subseteq n+1$. Then:
  1. (i)
    if $A=\{0,\ldots,n\}$ then $|\mathcal{C}_{A}(\vec{\kappa})|=2^{\kappa_{n}}$;
  2. (ii)
    if $A=\emptyset$ then $|\mathcal{C}_{A}(\vec{\kappa})|=\kappa_{n}$; and
  3. (iii)
    if $A$ and its complement $A^{c}$ are nonempty then C_A(κ)— = (κ_maxA^c)^κ_maxA.
Item (i) follows because $\mathcal{D}_{\{0,\ldots,n\}}$ is just the discrete topology on $\prod_{i\leq n}\kappa_{i}$. To show (ii), we induct on the length of $\vec{\kappa}$:
  • If $n=0$ then $\mathcal{C}_{\emptyset}(\vec{\kappa})$ is just the ring of continuous functions $\alpha(\kappa_{0})\to\mathbb{Z}_{2}$. These correspond to the finite-or-cofinite subsets of $\kappa_{0}$, of which there are $\kappa_{0}$ many.
  • Now suppose $|\mathcal{C}_{\emptyset}(\vec{\kappa})|=\kappa_{n}$ for any tuple $\vec{\kappa}$ of length $n$, and fix a continuous function $f:\prod_{i\leq n}\alpha(\kappa_{i})\to\mathbb{Z}_{2}$. Since the codomain of $f$ is discrete, the preimage under $f$ of $f(\infty,\ldots,\infty)$ is a neighborhood of $(\infty,\ldots,\infty)$. Such a neighborhood contains some rectangle $R=\prod_{i\leq n}J_{i}$, where $J_{i}\subseteq\alpha(\kappa_{i})$ is cofinite and contains $\infty$. We can then write the domain of $f$ as the disjoint union of finitely many subspaces, as follows:
    $\prod_{i\leq n}\alpha(\kappa_{i})$ $\cong\prod_{i\leq n}\left(J_{i}\amalg(\alpha(\kappa_{i})\setminus J_{i})\right)$
    $\cong\coprod_{A\subseteq n+1}\left(\prod_{i\in A}J_{i}\times\prod_{i\not\in A}(\alpha(\kappa_{i})\setminus J_{i})\right).$
    This is a disjoint union of $R$ and finitely many spaces of the form $\prod_{i\in A}J_{i}$, where $A\subseteq n+1$ is of size <n+1<n+1. Since $J_{i}$ is a cofinite subset of $\alpha(\kappa_{i})$ containing $\infty$, any space of this form is homeomorphic to $\mathcal{C}_{\emptyset}(\vec{\kappa}\restriction_{A})$, which by the inductive hypothesis has cardinality $\kappa_{\max A}$.
    We have thus described a bijection
    $\mathcal{C}_{\emptyset}(\vec{\kappa})\xrightarrow{\sim}\prod_{i\leq n}[\kappa_{i}]^{<\aleph_{0}}\times\{0,1\}\times\prod_{A\in[n+1]^{<n+1}}\kappa_{\max A},$
    where the first coordinates record the rectangle $R$ on which $f$ is constant, the second records this constant value, and the last coordinates record the restrictions of $f$ to each $\mathcal{C}_{\emptyset}(\vec{\kappa}\restriction_{A})$. The latter product has cardinality $\kappa_{n}$, so $|\mathcal{C}_{\emptyset}(\vec{\kappa})|=\kappa_{n}$ as desired to show (ii).
To conclude, we now need only note that by Fact 2 above we have
$\mathcal{C}_{A}(\vec{\kappa})\cong\mathcal{C}_{\emptyset}(\vec{\kappa}^{A})^{\prod\limits_{j\in A}\kappa_{j}}$
so that (iii) follows from (ii).  ∎
We will also need the following lemma; note that its statement leverages the abuse of notation from Fact 1.
Lemma 3.2.
For every tuple $\vec{\kappa}$ of length $n+1$,
$\left|\bigcap_{i\leq n}\mathcal{C}_{\{i\}}(\vec{\kappa})\right|\geq 2^{\kappa_{0}}.$
Fix $B\subseteq\kappa_{0}$. We define
$\Delta B$ $:\kappa_{0}\times\cdots\times\kappa_{n}\to\mathbb{Z}_{2},$
$\Delta B(\vec{x})$ $=\begin{cases}1&\text{if $x_{i}=x_{j}\in B$ for all $i,j$,}\\ 0&\text{otherwise.}\end{cases}$
To see $\Delta B\in\mathcal{C}_{\{i\}}$ for all $i$ it suffices to show it is constant in a neighborhood of any limit point in $X(\vec{\kappa})^{-}$. So fix $\vec{x}\in X(\vec{\kappa})^{-}$, and let $x_{j}$ be a coordinate which is not $\infty$. If $\vec{x}$ is the constant tuple $\langle x_{j},\ldots,x_{j}\rangle$, then $\vec{x}$ is isolated, since $x_{j}\neq\infty$; so certainly $\Delta B$ is constant on a neighborhood of $\vec{x}$. Otherwise, $\vec{x}$ lies in the open “punctured hyperplane”
$\{\vec{y}\in X(\vec{\kappa})^{-}:x_{j}=y_{j}\}\setminus\{\langle x_{j},\ldots,x_{j}\rangle\},$
on which $\Delta B$ is constant. [^2] Note that if $x_{j}\not\in\kappa_{0}$, the removal of the latter point is vacuous.
Lastly notice that $\Delta:2^{\kappa_{0}}\to\bigcap_{i\leq n}\mathcal{C}_{\{i\}}(\vec{\kappa})$ is certainly injective, since if $n\in B\triangle B^{\prime}$ then $\Delta B(n,\ldots,n)\neq\Delta B^{\prime}(n,\ldots,n)$.  ∎
We are now ready to prove Theorem A. We will benefit from the following notation: for a fixed tuple $\vec{\kappa}$, write
$\mathcal{P}=\prod_{i=0}^{n-k}\kappa_{i},\qquad\mathcal{Q}=\prod_{i=n-k+1}^{n}\kappa_{i}.$
We equip $\mathcal{P}$ and $\mathcal{Q}$ with the coordinatewise partial order, so that $|\mathcal{P}|=\kappa_{n-k}$ and $\operatorname{cf}\mathcal{Q}=\operatorname{cf}(\kappa_{n})$.
We also make the following definition for convenience:
Definition 3.3.
A subset $A\subseteq n+1$ is $k$-terminal if it contains the last $k$ elements of $n+1$, i.e., $\{n-k+1,\ldots,n\}\subseteq A$.
Proposition 3.4.
Suppose $\kappa_{n-k}<\operatorname{cf}(\kappa_{n})$, and let $\vec{f}=\langle f_{A}\rangle$ be a (k1)(k-1)-cochain. Then for any $k$-terminal $A\in[n+1]^{k+1}$ there is a cofinal subset $C\subseteq\mathcal{Q}$ such that the map
$C$ $\to\mathcal{C}_{\emptyset}(\kappa_{0},\ldots,\kappa_{n-k})$
$\vec{y}$ $\mapsto(d\vec{f})_{A}(-,\vec{y})$
is constant.
Expanding the definition of $d\vec{f}$ using that $A$ is $k$-terminal,
$(d\vec{f})_{A}=f_{\{n-k+1,\ldots,n\}}+\sum_{j=n-k+1}^{n}f_{A\setminus\{j\}}.$
The codomain of $f_{\{n-k+1,\ldots,n\}}(-,\vec{y}),$ i.e., $\mathcal{C}_{\emptyset}(\kappa_{0},\ldots,\kappa_{n-k}),$ has cardinality $\kappa_{n-k}<\operatorname{cf}\mathcal{Q}$, so the first summand is constant on a cofinal subset $C_{0}$ of $\mathcal{Q}$.
For the remaining summands, fix $j\in\{n-k+1,\ldots,n\}$. Since $f_{A\setminus\{j\}}\in\mathcal{C}_{A\setminus\{j\}}$, by Fact 2 we have that for any $\vec{x}\in\mathcal{P}$
$f_{A\setminus\{j\}}(\vec{x},-)\in\mathcal{C}_{\{n-k+1,\ldots,\hat{j},\ldots,n\}}(\kappa_{n-k+1},\ldots,\kappa_{n}).$
That is, if $\vec{y}\in X(\kappa_{n-k+1},\ldots,\kappa_{n})$ satisfies $y_{i}=\infty$ iff $i=j$, then $g_{j}(\vec{x},-)$ must be constant in a neighborhood of $\vec{y}$. Thus there is $\beta(\vec{x},j)\in\kappa_{j}$ such that $g_{j}(\vec{x},\vec{y})$ is independent of $y_{j}$ whenever $y_{j}>\beta(\vec{x},j)$. Taking
$\vec{\beta}(\vec{x})=\langle\beta(\vec{x},j):n-k+1\leq j\leq n\rangle\in\mathcal{Q},$
we find that $\sum_{j=n-k+1}^{n}g_{j}(\vec{x},\vec{y})$ is independent of $\vec{y}$ whenever $\vec{y}>\vec{\beta}(\vec{x})$ coordinatewise.
The above describes a map $\vec{x}\mapsto\vec{\beta}(\vec{x}):\mathcal{P}\to\mathcal{Q}$; so again by our assumption the image of this map must be bounded in $\mathcal{Q}$. Taking $\vec{\beta}$ to be a common upper bound in $\mathcal{Q}$ of all the $\vec{\beta}(\vec{x})$, we find that $\sum_{j=n-k+1}^{n}g_{j}(\vec{x},\vec{y})$ is constant for $\vec{y}>\vec{\beta}$.
To conclude, we let $C=\{\vec{y}\in C_{0}:\vec{y}>\vec{\beta}\}$. Then $C$ remains cofinal in $\mathcal{Q}$, and the above arguments and the expanded definition of $d\vec{f}$ show that $(d\vec{f})_{A}(-,\vec{x})$ is constant on $C$.  ∎
Proposition 3.4 describes a condition satisfied by any $k$-coboundary. As the final step, we construct a $k$-cocycle which does not satisfy this condition.
Proposition 3.5.
Suppose $2^{\kappa_{0}}\geq\kappa_{n-k+1}$. Then there is a $k$-cocycle $\vec{f}=\langle f_{A}\rangle$ such that for any $k$-terminal AA the map $\vec{y}\mapsto f_{A}(-,\vec{y})$ is not constant on any cofinal subset of $\mathcal{Q}$. In particular, if $\kappa_{n-k}<\operatorname{cf}(\kappa_{n})$, then $\vec{f}$ represents a nonzero cohomology class in $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$.
Starting from Lemma 3.2, we note that
$\left|\bigcap_{i\leq n-k}\mathcal{C}_{\{i\}}(\kappa_{0},\ldots,\kappa_{n-k})\right|$ $\geq 2^{\kappa_{0}}$
and the latter is at least $\kappa_{n-k+1}$ by our hypothesis. So fix an injection
$\sigma:\kappa_{n-k+1}\hookrightarrow\bigcap_{i\leq n-k}\mathcal{C}_{\{i\}}(\kappa_{0},\ldots,\kappa_{n-k}),$
and for $A\in[n+1]^{k+1}$ define
$f_{A}$ $:\mathcal{D}_{A}(\vec{\kappa})\to\mathbb{Z}_{2},$
$f_{A}(\vec{x})$ $:=\begin{cases}\sigma(x_{n-k+1})(x_{0},\ldots,x_{n-k})\quad&\text{if $A$ is $k$-terminal,}\\ 0\quad&\text{otherwise.}\end{cases}$
Notice first that each $f_{A}$ is well-defined, since if $A=\{i\}\cup\{n-k+1,\ldots,n\}$ and $\vec{x}\in\mathcal{D}_{A}(\vec{\kappa})$ then $x_{i},x_{n-k+1}\neq\infty$.
To see that each $f_{A}$ is continuous, assume $A$ is $k$-terminal; otherwise $f_{A}$ is identically zero and thus certainly continuous. Recalling the isomorphism from Fact 2
$\mathcal{C}_{\{i\}}(\kappa_{0},\ldots,\kappa_{n-k})^{Q}$ $\cong\mathcal{C}_{\{i\}\cup\{n-k+1,\ldots,n\}}(\vec{\kappa})$
$=\mathcal{C}_{A}(\vec{\kappa}),$
we see that $f_{A}$ arises from the function
$\langle x_{n-k+1},\ldots,x_{n}\rangle\mapsto\sigma(x_{n-k+1})$
and thus lies in $\mathcal{C}_{A}(\vec{\kappa})$.
To see $\vec{f}$ defines a cocycle, choose any $A^{\prime}\in[n+1]^{k+2}$. Then by definition
$(d\vec{f})_{A^{\prime}}=\sum_{i\in A^{\prime}}f_{A^{\prime}\setminus\{i\}}.$
We will show this sum always vanishes. Again, the case where $A^{\prime}$ is not $k$-terminal is immediate since every term is zero. So suppose $A^{\prime}=\{i_{0},i_{1}\}\cup\{n-k+1,\ldots,n\}$. Then the only nonzero summands are those indexed by $i=i_{0},i_{1}$, and $d\vec{f}$ is independent of these coordinates; so
$(df)_{A^{\prime}}(\vec{x})$ $=f_{A^{\prime}\setminus\{i_{0}\}}(\vec{x})+f_{A^{\prime}\setminus\{i_{1}\}}(\vec{x})$
$=0.$
So $\vec{f}$ is a cocycle, as desired.
We conclude by noting that, for any $k$-terminal $A\in[n+1]^{k+1}$, the map
$\mathcal{Q}$ $\to\mathcal{C}_{\emptyset}(\kappa_{0},\ldots,\kappa_{n-k})$
$\vec{x}$ $\mapsto f_{A}(-,\vec{x})$
$\equiv\sigma(x_{n-k+1})$
cannot be constant on a cofinal subset of $\mathcal{Q}$; any such subset certainly contains tuples whose first coordinates disagree, and the injectivity of $\sigma$ guarantees that the $f_{A}(-,\vec{x})$ will be distinct for tuples with differing first coordinate. Hence $\vec{f}$ is nontrivial, as desired.  ∎
4. Constructions via Fubini partitions
As mentioned above, our second main computational tool is the notion of a $k$-Fubini partition of $X(\vec{\kappa})^{-}$. To define this, we begin with a more careful dissection of the topology of $X(\vec{\kappa})$. We make the following notation:
Definition 4.1.
For any $A\subseteq n+1$, write $F_{A}$ for the AA-facet of $X(\vec{\kappa})$, i.e.,
$F_{A}=\{\vec{x}\in X(\vec{\kappa}):x_{i}\neq\infty\text{ if and only if }i\in A\}.$
Note that $X(\vec{\kappa})$ is the disjoint union of the facets $F_{A}$ for $A\subseteq n+1$, and for any $A\subseteq n+1$ we have
$\mathcal{D}_{A}(\vec{\kappa})=\bigcup_{A\subseteq B\subseteq n+1}F_{B}.$
A basis for the topology of $X(\vec{\kappa})$ can then be given as follows: for any $A\subseteq n+1$, any $\vec{x}\in F_{A}$, and any tuple of finite sets
$\vec{\mathcal{E}}\in\prod_{j\not\in A}[\kappa_{j}]^{<\aleph_{0}},$
we define the open set
$\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}}):=\{\vec{y}\in X(\vec{\kappa}):\vec{y}\upharpoonright_{A}=\vec{x}\upharpoonright_{A}\text{ and }y_{j}=\infty\text{ or }y_{j}\not\in\mathcal{E}_{j}\text{ for all }j\not\in A\}.$
That is, $\vec{y}\neq\vec{x}$ lies in this neighborhood if and only if it agrees with $\vec{x}$ wherever $x_{j}$ is $\infty$, and all other coordinates are either $\infty$ or lie outside the corresponding finite set $\mathcal{E}_{j}$. Figure 1 illustrates the sets $F_{A}$ and $\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})$ for a particular choice of $\vec{x}$ and $\vec{\mathcal{E}}$.
[…]
Figure 1. The facets $F_{A}$ and basic open set $\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})$. Here $\vec{\kappa}=\langle\aleph_{0},\aleph_{1},\aleph_{2}\rangle$ and $A=\{1\}$.
Now, we define the main object of this section.
Definition 4.2.
Let $\vec{\kappa}=\langle\kappa_{0},\ldots,\kappa_{n}\rangle$ be a tuple of cardinals, and fix $1\leq k\leq n$. We call an $(n+1)$-element partition $\langle Y_{i}:i\leq n\rangle$ of $X(\vec{\kappa})^{-}$ $k$-Fubini if, for any $A\in[n+1]^{\geq k+1}$, the following criteria hold:
  1. (1)
    Whenever $i\not\in A$ the set $Y_{i}$ is clopen in $\mathcal{D}_{A}$; and
  2. (2)
    for any $\vec{x}\in F_{A}$, there is $\vec{\mathcal{E}}\in\prod\limits_{j\not\in A}[\kappa_{j}]^{<\aleph_{0}}$ such that N_A(→x, →E) ⊆⋃_i ∈A Y_i.
Remark 4.3.
We make a few observations.
  • Condition (1) will allow us to define continuous functions $f_{A}$ on $\mathcal{D}_{A}$ by defining them piecewise on the $Y_{i}\cap\mathcal{D}_{A}$; continuity only needs to be checked on the individual pieces.
  • Condition (2) equivalently asserts that for any $A\in[n+1]^{\geq k+1}$, the set $\bigcup_{i\in A}Y_{i}$ contains an open neighborhood of $F_{A}$, i.e.
    $F_{A}\subseteq\text{int}\big(\bigcup_{i\in A}Y_{i}\big).$
    However, it is the more verbose condition (2) which will be most useful to us.
  • It is immediate from the definition that a $k$-Fubini partition is $\ell$-Fubini whenever $k\leq\ell$; moreover, if $\vec{\kappa}\neq\emptyset$ then (1)(-1)-Fubini partitions never exist, and any partition into clopen sets is $n$-Fubini. Consequently, for every nonempty tuple there is a $k_{0}$ such that a $k$-Fubini partition exists if and only if $k_{0}\leq k$.
  • Lastly, we observe that an (n1)(n-1)-Fubini partition is the same as a partition of the product $\prod_{i\leq n}\kappa_{i}$ into $n+1$ many sets $Y_{i}$ such that $Y_{i}$ is “finite in ii”– in the sense that for any $\vec{x}\in\prod_{j\neq i}\kappa_{i}$, there are only finitely many $\vec{y}\in Y_{i}$ extending $\vec{x}$. This is the sense in which the partition witnesses a failure of Fubini’s theorem for $X(\vec{\kappa})$.
We defer discussion of the (non-)existence of Fubini partitions to the end of the section, focusing instead on their uses.
We will actually use the following strengthening of condition (1), which uses condition (2):
Lemma 4.4.
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-Fubini partition of $X(\vec{\kappa})$. Then for any $A\in[n+1]^{\geq k+1}$ and any $i\not\in A$,
$\mathcal{D}_{A}\cap Y_{i}=\mathcal{D}_{A\cup\{i\}}\cap Y_{i}.$
Fix $A\in[n+1]^{\geq k+1}$ and $i\not\in A$. For any $B\supseteq A$ we have by Lemma 4.3 that $F_{B}\subseteq\cup_{i\in B}Y_{i}$. In particular, if $i\not\in B$ then $F_{B}$ and $Y_{i}$ are disjoint; that is,
$\coprod_{\begin{subarray}{c}B\supseteq A\\ i\not\in B\end{subarray}}F_{B}\cap Y_{i}=\emptyset.$
Adding the $F_{B}$ for $B\supseteq A,i\in B$ to both sides, we get
$\coprod_{B\supseteq A}F_{B}\cap Y_{i}=\coprod_{\begin{subarray}{c}B\supseteq A\\ i\in B\end{subarray}}F_{B}\cap Y_{i}.$
which becomes $\mathcal{D}_{A}\cap Y_{i}=\mathcal{D}_{A\cup\{i\}}\cap Y_{i}$ by the remark in Definition 4.1.  ∎
The purpose of Lemma 4.4 is to “upgrade” a continuous function $\mathcal{D}_{A\cup\{i\}}(\vec{\kappa})\to\mathbb{Z}_{2}$ to a continuous function $\mathcal{D}_{A}(\vec{\kappa})\to\mathbb{Z}_{2}$, when restricted to the subspace $\mathcal{D}_{A\cup\{i\}}\cap Y_{i}$. This, in tandem with condition (1), enables piecewise definitions of cocycles using functions which are not continuous on the whole of $\mathcal{D}_{A}$.
As a first example of this, we construct a trivializer of a $k$-cocycle using a $k$-Fubini partition:
Proposition 4.5.
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-Fubini partition of $X(\vec{\kappa})$, and suppose $\vec{f}\in\prod_{A\in[n+1]^{k+2}}\mathcal{C}_{A}(\vec{\kappa})$ is a tuple of continuous functions. For any $B\in[n+1]^{k+1}$, any $i\leq n$, and any $\vec{x}\in Y_{i}$, define
$g_{B}(\vec{x})=\begin{cases}f_{B\cup\{i\}}(\vec{x})&\text{if }i\not\in B,\\ 0&\text{otherwise.}\end{cases}$
Then $g_{B}\in\mathcal{C}_{B}(\vec{\kappa})$. Moreover, if $\vec{f}$ is a cocycle, then $d\vec{g}=\vec{f}$.
As remarked above, we need only check the continuity of $g_{B}$ restricted to each $\mathcal{D}_{B}\cap Y_{i}$. Continuity of $g_{B}$ on $\mathcal{D}_{B}\cap Y_{i}$ where $i\not\in B$ follows from the continuity of $f_{B\cup\{i\}}$ on $\mathcal{D}_{B\cup\{i\}}\cap Y_{i}$ and Lemma 4.4. If $i\in B$ then $g_{B}$ is constant on $\mathcal{D}_{B}\cap Y_{i}$ and hence continuous.
To see that $d\vec{g}=\vec{f}$, we first note that for any $A\in[n+1]^{k+2}$ we have
$(d\vec{g})_{A}=\sum_{i\in A}g_{A\setminus\{i\}}.$
Fix $j\leq n$ and $\vec{x}\in Y_{j}$. If $j\in A$ then $g_{A\setminus\{i\}}(\vec{x})=0$ unless $i=j$, in which case it is $f_{A}(\vec{x})$. If $j\not\in A$ then the sum is $\sum_{i\in A}f_{A\setminus\{i\}\cup\{j\}}(\vec{x})$ which is equal to $f_{A}(\vec{x})$ by the cocycle property of $\vec{f}$. Thus $d\vec{g}=\vec{f}$ as desired.  ∎
Corollary 4.6.
If there is a $k$-Fubini partition of $X(\vec{\kappa})^{-}$ then $H^{\ell}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$ for all $\ell\geq k$.
Figure 2 shows a table view of the piecewise definition of the $g_{B}$ for the case $k=1,n=3$.
$Y_{0}$ $Y_{1}$ $Y_{2}$ $Y_{3}$
$g_{01}$ 0 0 $f_{012}$ $f_{013}$
$g_{02}$ 0 $f_{012}$ 0 $f_{023}$
$g_{03}$ 0 $f_{013}$ $f_{023}$ 0
$g_{12}$ $f_{012}$ 0 0 $f_{123}$
$g_{13}$ $f_{013}$ 0 $f_{123}$ 0
$g_{23}$ $f_{023}$ $f_{123}$ 0 0
Figure 2. Given a 2-cocycle $\vec{f}=\langle f_{012},f_{013},f_{023},f_{123}\rangle$ and a $k$-Fubini partition $\langle Y_{0},Y_{1},Y_{2},Y_{3}\rangle$, the above piecewise definition of $\langle g_{B}:B\in[4]^{2}\rangle$ defines a trivializer of $\vec{f}$. (For brevity the columns are labeled $Y_{i}$ rather than $\mathcal{D}_{A}\cap Y_{i}$.)
We next use a $k$-Fubini partition of $X(\vec{\kappa})^{-}$ to construct an embedding of the group $H^{k}(X(\vec{\lambda})^{-},\underline{\mathbb{Z}_{2}})$ into $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ when $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise. A priori, one might try to construct an embedding by treating large enough coordinates as infinity and mapping unrestricted values to zero, as follows:
Definition 4.7.
For $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise and a function $f_{A}\in\mathcal{C}_{A}(\vec{\lambda})$, define the naive extension of $f_{A}$ by
[…]
where $\vec{x}^{\infty}$ is obtained from $\vec{x}$ by setting all $x_{i}>=\lambda_{i}$ equal to $\infty$.
This extends to a map on cochains $\vec{f}\mapsto\vec{f}^{\infty}$ by acting on the individual functions.
Figure 3 shows an example of the naive extension of a cochain from $X(\vec{\lambda})^{-}$ to $X(\vec{\kappa})^{-}$ for $\lambda\leq\kappa$.
$0\scriptstyle{(\aleph_{0})}$$1\scriptstyle{(\aleph_{0})}$$2\scriptstyle{(\aleph_{0})}$
$0\scriptstyle{(\aleph_{0})}$$1\scriptstyle{(\aleph_{0})}$$2\scriptstyle{(\aleph_{1})}$$\aleph_{0}$
$0\scriptstyle{(\aleph_{0})}$$1\scriptstyle{(\aleph_{0})}$$2\scriptstyle{(\aleph_{0})}$
$0\scriptstyle{(\aleph_{0})}$$1\scriptstyle{(\aleph_{0})}$$2\scriptstyle{(\aleph_{1})}$$\aleph_{0}$0
$0\scriptstyle{(\aleph_{0})}$$1\scriptstyle{(\aleph_{0})}$$2\scriptstyle{(\aleph_{0})}$
$0\scriptstyle{(\aleph_{0})}$$1\scriptstyle{(\aleph_{0})}$$2\scriptstyle{(\aleph_{1})}$$\aleph_{0}$0
Figure 3. The naive extension of a 1-cochain $\langle f_{01},f_{02},f_{12}\rangle$ defined on $X(\aleph_{0},\aleph_{0},\aleph_{0})^{-}$ (left) to $\langle f_{01}^{\infty},f_{02}^{\infty},f_{12}^{\infty}\rangle$ defined on $X(\aleph_{0},\aleph_{0},\aleph_{1})^{-}$ (right). Notice that the latter is not necessarily a cocycle.
This definition indeed maps cochains to cochains, and in the case $k=n$ it maps cocycles to cocycles. Since a restriction of a trivializer for $\vec{f}^{\infty}$ to $X(\lambda)^{-}$ yields a trivializer for $\vec{f}$, it follows that for top cohomology
$H^{n}(X(\vec{\lambda})^{-},\underline{\mathbb{Z}_{2}})\hookrightarrow H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\quad\text{whenever}\quad\vec{\lambda}\leq\vec{\kappa}.$
However, the naive extension fails to generate embeddings in the case of intermediate cohomology:
Example 4.8.
Let $\vec{\lambda}=\langle\aleph_{0},\aleph_{0},\aleph_{0}\rangle,\vec{\kappa}=\langle\aleph_{0},\aleph_{0},\aleph_{1}\rangle,$ and let $\vec{f}\in H^{1}(X(\vec{\lambda})^{-},\underline{\mathbb{Z}_{2}})$ be defined by
$f_{01}(\vec{x})$ $=f_{02}(\vec{x})=1,$
$f_{12}(\vec{x})$ $=0.$
The naive extension maps $f_{01}$ and $f_{12}$ to the constant-1 function and constant-0 function, respectively, on $X(\vec{\kappa})$, but $f_{02}^{\infty}(\vec{x})=0$ when $x_{2}>\aleph_{0}$. Thus the sum $f_{01}^{\infty}+f_{02}^{\infty}+f_{12}^{\infty}$ is not zero, so $\vec{f}^{\infty}$ is not a cocycle.
Thus in order to construct an embedding, we will modify the cochain $\vec{f}^{\infty}$ to a cocycle. In fact we will show that, given the existence of a $k$-Fubini partition, any $k$-cochain can be modified to a $k$-cocycle without affecting its values at any limit points.
Definition 4.9.
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-Fubini partition of $X(\vec{\kappa})^{-}$, and let $\vec{f}\in H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ be a $k$-cochain. Define the modification of $\vec{f}$ as follows: for any $A\in[n+1]^{k+1}$ and any $\vec{x}\in\mathcal{D}_{A}\cap Y_{i}$, define
$f^{\operatorname{mod}}_{A}(\vec{x})=\begin{cases}f_{A}(\vec{x})&\text{if }i\in A,\\ \sum_{j\in A}f_{A\setminus\{j\}\cup\{i\}}(\vec{x})&\text{if }i\not\in A.\end{cases}$
Lemma 4.10.
For any cochain $\vec{f}$, the modified tuple $\vec{f}^{\operatorname{mod}}$ is a cocycle. Moreover, for any $A\in[n+1]^{k+1}$ and any limit point $\vec{x}\in\mathcal{D}_{A}$, $f_{A}(\vec{x})=f_{A}^{\operatorname{mod}}(\vec{x})$.
We first show that each $f^{\operatorname{mod}}_{A}$ is continuous; by Remark 4.3 it suffices to show $f^{\operatorname{mod}}_{A}$ is continuous on each $\mathcal{D}_{A}\cap Y_{i}$. So let $\vec{x}\in\mathcal{D}_{A}\cap Y_{i}$:
  • If $i\in A$, let $U$ be a neighborhood of $\vec{x}$ in $Y_{i}$ such that $f_{A}$ is constant on $U$. Then $f^{\operatorname{mod}}_{A}$ is constant on $U\cap Y_{i}$ and hence continuous at $\vec{x}$.
  • If $i\not\in A$, then by Lemma 4.4 we have for each $j\in A$
    $\mathcal{D}_{A\setminus\{j\}}\cap Y_{i}=\mathcal{D}_{A\setminus\{j\}\cup\{i\}}\cap Y_{i}.$
    Since $f_{A\setminus\{j\}\cup\{i\}}$ is continuous on $\mathcal{D}_{A\setminus\{j\}\cup\{i\}}\cap Y_{i}$, there is a neighborhood $U_{j}\subseteq\mathcal{D}_{A\setminus\{j\}}\cap Y_{i}$ such that $f_{A\setminus\{j\}\cup\{i\}}$ is constant on $U_{j}$. Let $U=\mathcal{D}_{A}\cap\bigcap_{j\in A}U_{j}$; then in $U$ we have that each $f_{A\setminus\{j\}\cup\{i\}}$ is constant, so that
    $g_{A}=\sum_{j\in A}f_{A\setminus\{j\}\cup\{i\}}$
    is constant on $U$.
So $\vec{f}^{\operatorname{mod}}$ is a cochain. To check the cocycle condition, assume $k\leq n-1$ (if $k=n$ then the cocycle condition is automatic) and choose $B\in[n+1]^{k+2}$. Again take $\vec{x}\in Y_{i}$ and case split on whether $i\in B$:
  • If $i\in B$, then
    $\sum_{j\in B}f^{\operatorname{mod}}_{B\setminus\{j\}}(\vec{x})$ $=f^{\operatorname{mod}}_{B\setminus\{i\}}(\vec{x})+\sum_{j\in B\setminus\{i\}}f^{\operatorname{mod}}_{B\setminus\{j\}}(\vec{x})$
    $=\sum_{j\in B\setminus\{i\}}f_{B\setminus\{j\}}(\vec{x})+\sum_{j\in B\setminus\{i\}}f_{B\setminus\{j\}}(\vec{x})$
    $=0;$
  • If $i\not\in B$ then
    $\sum_{j\in B}f^{\operatorname{mod}}_{B\setminus\{j\}}(\vec{x})$ $=\sum_{j\in B}\sum_{k\in B}f_{B\setminus\{j,k\}\cup\{i\}}$
    $=0.$
Lastly, to see that modification preserves values at limit points, let $\vec{x}\in\mathcal{D}_{A}$ be a limit point, i.e., $\vec{x}\in F_{B}$ for some $A\subseteq B\subsetneq n+1$. Since $|B|\geq k+1$, condition (2) of the $k$-Fubini partition implies the existence of a neighborhood of $\vec{x}$ in which every point lies in $Y_{i}$ for some $i\in A$. Thus, in this neighborhood $f^{\operatorname{mod}}_{A}\equiv f_{A}.$  ∎
Figure 4 shows a table view of the piecewise definition of $\vec{f}^{\operatorname{mod}}$ for the case $k=1,n=3$.
$Y_{0}$ $Y_{1}$ $Y_{2}$ $Y_{3}$
$f^{\operatorname{mod}}_{01}$ $f_{01}$ $f_{01}$ $f_{01}+f_{02}$ $f_{03}+f_{13}$
$f^{\operatorname{mod}}_{02}$ $f_{02}$ $f_{01}+f_{12}$ $f_{02}$ $f_{03}+f_{23}$
$f^{\operatorname{mod}}_{03}$ $f_{03}$ $f_{01}+f_{13}$ $f_{02}+f_{23}$ $f_{03}$
$f^{\operatorname{mod}}_{12}$ $f_{01}+f_{02}$ $f_{12}$ $f_{12}$ $f_{13}+f_{23}$
$f^{\operatorname{mod}}_{13}$ $f_{01}+f_{03}$ $f_{13}$ $f_{12}+f_{23}$ $f_{13}$
$f^{\operatorname{mod}}_{23}$ $f_{02}+f_{03}$ $f_{12}+f_{13}$ $f_{23}$ $f_{23}$
Figure 4. Given a 1-cochain $\vec{f}=\langle f_{01},f_{02},f_{03},f_{12},f_{13},f_{23}\rangle$ and a $k$-Fubini partition $\langle Y_{0},Y_{1},Y_{2},Y_{3}\rangle$, the above piecewise definition of $\langle f^{\operatorname{mod}}_{A}:A\in[4]^{2}\rangle$ modifies $\vec{f}$ to a cocycle without affecting its values at any limit points. (Again, for brevity the columns are labeled $Y_{i}$ rather than $\mathcal{D}_{A}\cap Y_{i}$.)
Proposition 4.11.
Suppose $X(\vec{\kappa})^{-}$ admits a $k$-Fubini partition. Then for any $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise, there is an embedding
$H^{k}(X(\vec{\lambda})^{-},\underline{\mathbb{Z}_{2}})\hookrightarrow H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}}).$
Given a $k$-cocycle $\vec{f}$ defined on $X(\vec{\lambda})^{-}$, perform the naive extension to obtain a cochain on $X(\vec{\kappa})^{-}$, then take its modification to obtain a cocycle. If this cocycle were trivial, then by Lemma 4.10 the restriction of its trivializer to $X(\vec{\lambda})$ would trivialize $\vec{f}$ as well.  ∎
It now remains to determine when Fubini partitions exist. As we saw above, the interesting cases are $0\leq k<n$. We first consider when k<n1k<n-1.
Proposition 4.12.
Let $\vec{\kappa}$ be a weakly increasing tuple of cardinals, $0\leq k<n-1$. Then a $k$-Fubini partition of $X(\vec{\kappa})^{-}$ exists if and only if $\kappa_{n-k}=\aleph_{0}$, that is, the first nk+1n-k+1 entries of $\vec{\kappa}$ are all $\aleph_{0}$.
For the forward direction, suppose $\vec{\kappa}$ admits a $k$-Fubini partition. Then we simultaneously have that (i) $\vec{\kappa}$ also admits a $(k+1)$-Fubini partition (Remark 4.3), and that (ii) $H^{k+1}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$ (Corollary 4.6). The only way to reconcile these with Proposition 4.11 is if no $\vec{\lambda}\leq\vec{\kappa}$ has nonzero $(k+1)$-cohomology. But now consider the tuple
$\vec{\lambda}=\langle\underbrace{\aleph_{0},\aleph_{0},\ldots,\aleph_{0}}_{n-k},\underbrace{\aleph_{1},\aleph_{1},\ldots,\aleph_{1}}_{k+1}\rangle.$
By Theorem A, $\vec{\lambda}$ has nonzero $(k+1)$-cohomology (note $k+1<n$ by the assumption k<n1k<n-1). So it cannot lie below $\vec{\kappa}$ coordinatewise, which is equivalent to $\kappa_{n-k}=\aleph_{0}$.
For the backward direction, suppose $\kappa_{n-k}=\aleph_{0}$, and for $i\leq n$ set
$Y_{i}=\{\vec{x}\in X(\vec{\kappa}):\min\vec{x}=x_{i}\},$
where $\infty$ is treated as larger than any ordinal. [^3] Ties can be broken by an arbitrary scheme; for concreteness, we can prefer earlier coordinates, but we will never use this. These sets are disjoint, and since any tuple containing a non-$\infty$ coordinate has a minimum, they cover $X(\vec{\kappa})^{-}$. It remains to check conditions (1),(2); for this fix $A\in[n+1]^{\geq k+1}$. Note that there are only $k$ coordinates in the set $\{n-k+1,n-k+2,\ldots,n-1,n\}$, and $|A|>k$; thus there must be $i_{0}\in A$ such that $i_{0}\leq n-k$, and so by hypothesis $\kappa_{i_{0}}=\aleph_{0}$. So if $\vec{x}\in\mathcal{D}_{A}$, $\min\vec{x}<\aleph_{0}$. This lets us quickly verify (1): namely, for $j$ such that $x_{j}\neq\infty$, set
$\mathcal{E}_{j}=\{0,1,\ldots,\min\vec{x}\}\in[\kappa_{j}]^{<\aleph_{0}}.$
Then $\mathcal{N}(\vec{x},\vec{\mathcal{E}})$ is a neighborhood of $\vec{x}$, all of whose points have the same minimum. If $\vec{x}\in Y_{i}$ then this neighborhood lies within $Y_{i}$ and if $\vec{x}\not\in Y_{i}$ it lies outside $Y_{i}$. So $Y_{i}$ is clopen in $\mathcal{D}_{A}$. A similar approach shows (2): fix $\vec{x}\in F_{A}$ and for $j\not\in A$ set $\mathcal{E}_{j}=\{0,1,\ldots,x_{i_{0}}\}$. If $\vec{y}\in\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})$, then for every $j\not\in A$ we have $y_{j}>x_{i_{0}}=y_{i_{0}}$; so $\min\vec{y}\neq y_{j}$ and thus $\vec{y}\not\in Y_{j}$. We conclude that the minimum of $\vec{y}$ is among the $y_{i}$ for $i\in A$, i.e.,
$\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})\subseteq\bigcup_{i\in A}Y_{i},$
completing the proof of (2).  ∎
The remaining case is k=n1k=n-1. Here the backward direction of Proposition 4.12 still holds: if $\kappa_{1}=\aleph_{0}$ then the above defines an (n1)(n-1)-Fubini partition, thus trivializing the top cohomology of $X(\vec{\kappa})^{-}$. On the other hand, Aoki showed that if $\kappa_{i}\geq\aleph_{i}$ for all $i$, then $H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$. Our converse to Aoki’s result comes from the requisite strengthening of the above forward direction for k=n1k=n-1:
Proposition 4.13.
Suppose there is $i_{0}\leq n$ such that $\kappa_{i_{0}}<\aleph_{i_{0}}$. Then an (n1)(n-1)-Fubini partition exists for $\vec{\kappa}$.
We will need a few ingredients for the proof.
  • First: we fix once and for all a family of injections
    $\langle e_{\beta}:\beta\hookrightarrow|\beta|\bigm|\beta<\kappa_{n}\rangle.$
    For a subset $A\subseteq n+1$ of size $\geq 2$ and a point $\vec{x}\in\prod_{i\in A}\kappa_{i}$, we define the virtual minimum $\operatorname{vmin}\vec{x}$ of $\vec{x}$ (with respect to the $e_{\beta}$) inductively on the size of $A$, as follows:
    • If $|A|=2$, then $\operatorname{vmin}\vec{x}=\min\vec{x}$ (again, breaking ties arbitrarily).
    • If $|A|\geq 3$, then the virtual minimum of $\vec{x}$ is defined as
      $\operatorname{vmin}\vec{x}:=\operatorname{vmin}\langle e_{\max\vec{x}}(x_{i}):i\in A,i\neq i_{\max}\rangle,$
      where $i_{\max}$ is the index of $\max{}\vec{x}$.
    Algorithmically: remove the maximum element from the tuple, apply its corresponding injection to the remaining elements, then repeat until the tuple consists of just two elements, at which point the virtual minimum is just the minimum.
    Certainly this process terminates, and designates exactly one element of the tuple as its virtual minimum. We denote the index of $\operatorname{vmin}\vec{x}$ by $\iota_{\operatorname{vmin}}(\vec{x})$.
  • Second: as noted in Remark 4.3, in the case k=n1k=n-1, conditions (1) and (2) of Definition 4.2 reduce to the single condition
    $(3)$ For every $j\leq n$ and every $\vec{x}^{j}=\langle x_{i}\in\kappa_{i}:i\neq j\rangle$, there are only finitely many possible values of $x_{j}\in\kappa_{j}$ such that the resulting tuple $\vec{x}$ lies in $Y_{j}$.
    So we will need only construct a partition $\langle Y_{i}\rangle$ satisfying (3).
  • Third: we in fact can assume $i_{0}=n$, i.e., $\kappa_{n}<\aleph_{n}$. This is because of the general fact that an $(i_{0}-1)$-Fubini partition of $X(\kappa_{0},\ldots,\kappa_{i_{0}})^{-}$ yields an (n1)(n-1)-Fubini partition of $X(\vec{\kappa})^{-}$ by setting the remaining $Y_{i}$ equal to the empty set.
With these in hand, we are ready to prove Proposition 4.13.
Assume $\kappa_{n}<\aleph_{n}$. We first claim that $\kappa_{n}<\aleph_{n}$ (i.e., $i_{0}=n$ in the notation of the statement) implies $\operatorname{vmin}\vec{x}<\aleph_{0}$. We prove this by induction on the length $n$:
  • If $n=1$ then $\vec{\kappa}=\langle\aleph_{0},\aleph_{0}\rangle$ and the claim is immediate.
  • Suppose the claim holds for tuples of length $n$, and let $\vec{\kappa}=\langle\kappa_{i}:i\leq n+1\rangle$ with $\kappa_{n+1}<\aleph_{n+1}$. Then, given $\vec{x}\in\prod_{i\leq n+1}\kappa_{i}$, we have $|\max\vec{x}|<\kappa_{n+1}$, so that for each $i\neq i_{\max}$ we have
    $e_{\max\vec{x}}(x_{i})<|\max\vec{x}\,|<\kappa_{n+1}<\aleph_{n+1}.$
    Therefore $\langle e_{\max\vec{x}}(x_{i}):i\neq i_{\max}\rangle$ is a tuple of length $n$ whose maximum is below $\aleph_{n}$; the inductive hypothesis then implies its virtual minimum, and thus the virtual minimum of $\vec{x}$, is finite.
Now, for $j\leq n$ define
$Y_{j}:=\left\{\vec{x}\in\prod_{i\leq n}\kappa_{i}:\iota_{\operatorname{vmin}}(\vec{x})=j\right\}.$
We claim that if the virtual minimum of $\vec{x}$ is finite, then the $Y_{j}$ satisfy (3). We again prove this by induction on the length of the tuple:
  • If $n=1$ then again $\vec{\kappa}=\langle\aleph_{0},\aleph_{0}\rangle$. Fixing one of these two coordinates, we see that only finitely many choices of the other coordinate enable it to be the (virtual) minimum of $\vec{x}$.
  • Now suppose the claim holds for tuples of length $n$, let $\vec{\kappa}=\langle\kappa_{0},\ldots,\kappa_{n+1}\rangle$ have finite virtual minimum, and fix all $x_{i}\in\kappa_{i}$ except a single coordinate $x_{j}$. Suppose we have chosen $x_{j}$ such that $\iota_{\operatorname{vmin}}(\vec{x})=j$. Then $j$ is also the virtual minimal index of $\langle e_{\max\vec{x}}(x_{i}):i\neq i_{\max}\rangle$, a tuple of length $n$ with finite virtual minimum. Thus by induction $e_{\max\vec{x}}(x_{j})$ comes from a finite set of possibilities, which pulls back to a finite set of possibilities for $x_{j}$ since the $e_{\beta}$ are injective.
So the $\langle Y_{j}\rangle$ form an (n1)(n-1)-Fubini partition.  ∎
Corollary 4.14.
Let $\vec{\kappa}$ be a tuple of cardinals such that $\kappa_{i}<i$ for some $i\leq n$. Then $H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$.
By Corollary 4.6 and Proposition 4.13.  ∎
5. Cocycles as pointwise limits of coboundaries
In this section we prove Theorem C. In the process, we will describe a technique for reducing questions about intermediate cohomology $(0<k<n)$ to questions about top cohomology $(k=n)$. We first note the existence of a certain restriction map between Čech complexes, which is well-defined on cohomology:
Definition 5.1.
Let $\vec{\kappa}$ be a tuple of cardinals of length $n+1$, $0<k<n$. Fix also some $B\in[n+1]^{k+1}$. The face restriction map
$\operatorname{Res}^{k}_{B}:\check{C}(\vec{\kappa})\to\check{C}^{k}(\vec{\kappa}\upharpoonright_{B})$
is defined as follows: for $\vec{x}\in\prod_{i\in B}\kappa_{i},$ extend $\vec{x}$ to $\vec{x}^{\infty}$ by
$x^{\infty}_{i}=\begin{cases}x_{i},&i\not\in B,\\ \infty,&i\in B.\end{cases}$
Then, given a tuple $\vec{f}=\langle f_{A}:A\in[n+1]^{k+1}\rangle\in\check{C}^{k}(\vec{\kappa})$, $\operatorname{Res}^{k}_{B}\vec{f}$ is the single map $f_{B}^{\infty}$ which sends $\vec{x}$ to $f_{B}(\vec{x}^{\infty})$.
Proposition 5.2.
For $0<k<n$ and $B\in[n+1]^{k+1}$, $\operatorname{Res}^{k}_{B}$ is well-defined on cohomology.
The assertion that $\operatorname{Res}^{k}_{B}$ maps cocycles to cocycles is nearly vacuous, since top cocycles are just ordinary set functions. To see that coboundaries are mapped to coboundaries, suppose $\vec{f}=d\vec{g}$ for some $\vec{g}\in\check{C}^{k-1}(\vec{\kappa})$. For each $j\in B$ define $g_{B\setminus\{j\}}^{\infty}$ exactly as for $\vec{f}$, so that $\vec{x}\in\prod_{i\in B}\kappa_{i}$ is mapped to $g_{B\setminus\{j\}}(\vec{x}^{\infty})$. Now each $g_{B\setminus\{j\}}^{\infty}\in\check{C}_{B\setminus\{j\}}(\vec{\kappa}\upharpoonright_{B})$, and $\sum_{j\in B}g_{B\setminus\{j\}}^{\infty}=f_{B}^{\infty}=\operatorname{Res}^{k}_{B}(\vec{f})$.  ∎
As an immediate corollary, we note that if $\operatorname{Res}^{k}_{B}(\vec{f})$ is nonzero in the top cohomology of $X(\vec{\kappa}\upharpoonright_{B})^{-}$, then $\vec{f}$ must be nonzero in $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$. This is the manner in which we use the face restriction map to reduce questions on intermediate cohomology to questions of top cohomology.
We next examine the cocycle condition more closely.
Definition 5.3.
Fix $\vec{\kappa}=\langle\kappa_{0},\ldots,\kappa_{n}\rangle,B\in[n+1]^{k+1}$, and let $\vec{f}\in H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ be a $k$-cocycle. For a tuple
$\vec{x}=\langle x_{i}\in\kappa_{i}\cup\{\infty\}:i\not\in B\rangle\in X(\vec{\kappa}^{B}),$
and a choice of remaining coordinates $\vec{y}\in\prod_{i\in B}\kappa_{i}$, we set
$\varphi_{\vec{f},B,\vec{x}}(\vec{y})=f_{B}(\vec{x},\vec{y}).$
These choices assemble into a function
$\vec{x}\mapsto\varphi_{\vec{f},B,\vec{x}}:X(\vec{\kappa}^{B})\to Z^{k}(\vec{\kappa}\upharpoonright_{B})$
which we denote by $\varphi_{\vec{f},B}$.
Note that the codomain of $\varphi_{\vec{f},B}$ is naturally equipped with the product topology (i.e., that of pointwise convergence).
Proposition 5.4.
For any $k$-cocycle $\vec{f}$, $B$, the map $\varphi_{\vec{f},B}$ is continuous. Moreover, for any $\vec{x}\in\prod_{i\not\in B}\alpha(\kappa_{i})$ except possibly $(\infty,\ldots,\infty)$, $\varphi_{\vec{f},B}(\vec{x})$ is a coboundary.
First notice that $Z^{k}(\vec{\kappa}\upharpoonright_{B})=\{f:\prod_{i\in B}\kappa_{i}\to\mathbb{Z}_{2}\}$ is homeomorphic to a product of one copy of $\mathbb{Z}_{2}$ for each $\vec{y}\in\prod_{i\in B}\kappa_{i}$; thus to show $\varphi_{\vec{f},B}$ is continuous, it suffices to show each of the coordinate functions
$\varphi_{\vec{f},B}^{\vec{y}}\quad$ $:\quad X(\vec{\kappa}^{B})\to\mathbb{Z}_{2}$
$\vec{x}\in\prod_{i\not\in B}\alpha(\kappa_{i})\quad$ $\mapsto\quad f_{B}(\vec{x},\vec{y})$
is continuous. This in turn, holds if and only if the support of $\varphi_{\vec{f},B}^{\vec{y}}$ is a clopen subset of $\prod_{i\not\in B}\alpha(\kappa_{i})$. But continuity of $f_{B}$ guarantees that
$\operatorname{supp}f_{B}\subseteq\left(\prod_{i\not\in B}\alpha(\kappa_{i})\right)^{\prod_{i\in B}\kappa_{i}}$
is clopen; since $\operatorname{supp}(\varphi_{\vec{f},B}^{\vec{y}})$ is exactly the projection of this set onto the $\vec{y}$-th coordinate, and $\prod_{i\not\in B}\alpha(\kappa_{i})$ is compact, the projection must be clopen as well. So $\varphi_{\vec{f},B}$ is continuous.
Now, let $\vec{x}\in\prod_{i\not\in B}\alpha(\kappa_{i})$ be a point other than $(\infty,\ldots,\infty)$, so there exists $j\not\in B$ such that $x_{j}\neq\infty$. Set $A=B\cup\{j\}$, and note that the cocycle condition for $\vec{f}$ guarantees that
$f_{B}=\sum_{i\in B}f_{B\setminus\{i\}\cup\{j\}}.$
So we obtain
$\varphi_{\vec{f},B}(\vec{x})=f_{B}(\vec{x},-)=\sum_{i\in B}f_{B\setminus\{i\}\cup\{j\}}(\vec{x},-)$
and the latter expresses $\varphi_{\vec{f},B}(\vec{x})$ as a sum of functions, each mod-finite constant in one of the coordinates $i\in B$.  ∎
Proposition 5.4 can be construed as stating that a cocycle $\vec{f}$ can be written as a certain pointwise limit of coboundaries. The converse also holds:
Proposition 5.5.
Let $\varphi:X(\vec{\kappa}^{B})\to Z^{k}(\vec{\kappa}\upharpoonright_{B})$ be a continuous map which sends all points except possibly $(\infty,\ldots,\infty)$ to coboundaries. Then there exists a $k$-cocycle $\vec{f}\in H^{k}(\vec{\kappa},\underline{\mathbb{Z}_{2}})$ such that $\varphi(\infty,\ldots,\infty)=\operatorname{Res}^{k}_{B}(\vec{f})$.
Let $A\in[n+1]^{k+1},\vec{x}\in\mathcal{D}_{A}$. To define $f_{A}$, we distinguish three cases:
  • If $A=B$, we set $f_{A}(\vec{x})=\varphi(\vec{x}^{B})(\vec{x}\upharpoonright_{B})$.
  • If $A=B\setminus\{s\}\cup\{t\}$, where $s\in B,t\not\in B$, then by assumption $\vec{x}^{B}\in X(\vec{\kappa}^{B})^{-}$. Thus $\varphi(\vec{x}^{B})\in Z^{k}(\vec{\kappa}\upharpoonright_{B})$ is a coboundary: there exist functions
    $\langle g_{B\setminus\{i\}}\in\mathcal{C}_{B\setminus\{i\}}(\vec{\kappa}\upharpoonright_{B}):i\in B\rangle$
    such that $\varphi(\vec{x}^{B})=\sum_{i\in B}g_{B\setminus\{i\}}$. We set $f_{A}(\vec{x}):=g_{B\setminus\{s\}}(\vec{x}\upharpoonright_{B})$. Note this definition depends only on $\vec{x}^{B}$ and $s$; in particular it is independent of $t$.
  • In all other cases we set $f_{A}$ to be identically zero.
We first show that the $f_{A}$ so defined form a cochain, i.e., are each continuous. If $A=B$, then continuity of $f_{A}$ is exactly continuity of $\varphi$, and if $|A\triangle B|>2$ then $f_{A}$ is zero, hence certainly continuous. So suppose $A=B\setminus\{s\}\cup\{t\},s\in B,t\not\in B$. In this case $f_{A}(\vec{x})=g_{B\setminus\{s\}}(\vec{x}\upharpoonright_{B})$, where $g_{B\setminus\{s\}}\in\mathcal{C}_{B\setminus\{s\}}(\vec{\kappa}\upharpoonright_{B})$. As noted above, this definition is independent of $t$, so that $g_{B\setminus\{s\}}$ remains continuous on $\mathcal{D}_{B\setminus\{s\}\cup\{t\}}$. This shows $\vec{f}\in\check{C}^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$.
To see that $\vec{f}$ is moreover a $k$-cocycle, let $A^{\prime}\in[n+1]^{k+2}$. First suppose $B\subseteq A^{\prime},$ so that $A^{\prime}=B\cup\{t\}$ for some $t\not\in B$. We calculate, for any $\vec{x}\in\mathcal{D}_{A^{\prime}}(\vec{\kappa})$,
$\sum_{i\in A^{\prime}}f_{A^{\prime}\setminus\{i\}}(\vec{x})$ $=f_{B}(\vec{x})+\sum_{s\in B}f_{B\setminus\{s\}\cup\{t\}}(\vec{x})$
$=f_{B}(\vec{x})+\sum_{s\in B}g_{B\setminus\{s\}}(\vec{x}\upharpoonright_{B})$
$=\varphi(\vec{x}^{B})(\vec{x}\upharpoonright_{B})+\varphi(\vec{x}^{B})(\vec{x}\upharpoonright_{B})$
$=0,$
as required. If instead $A^{\prime}$ contains exactly two elements not in $B$, write $A^{\prime}=B\setminus\{s_{0}\}\cup\{t_{0},t_{1}\}$ where $s_{0}\in B$ and $t_{0},t_{1}\not\in B$ are distinct. Then we calculate
$\sum_{i\in A^{\prime}}f_{A^{\prime}\setminus\{i\}}(\vec{x})$ $=f_{B\setminus\{s_{0}\}\cup\{t_{0}\}}(\vec{x})+f_{B\setminus\{s_{0}\}\cup\{t_{1}\}}(\vec{x})$
$+\sum_{s_{1}\in B\setminus\{s_{0}\}}f_{B\setminus\{s_{0},s_{1}\}\cup\{t_{0},t_{1}\}}(\vec{x}).$
The first two summands are equal due to the independence of the definition from $t_{0},t_{1}$; in the latter sum all functions are identically zero. So the overall sum is zero.
Lastly, if $A^{\prime}$ contains at least three elements not in $B$, then each $A^{\prime}\setminus\{i\}$ contains at least two elements not in $B$, so that each term in the sum is zero. This concludes the proof that $\vec{f}$ is a cocycle.
We finish by observing that substituting $(\infty,\ldots,\infty)$ in the definition of $\varphi_{\vec{f},B}$ yields a map sending $\vec{x}$ to $f_{B}(\vec{x}^{\infty})$, which is exactly $\operatorname{Res}^{k}_{B}(\vec{f})$.  ∎
This, along with Proposition 5.5, completes the proof of Theorem C.
To finish this section, we apply Theorem C to settle a case not handled by the methods of sections 3 or 4:
Proposition 5.6.
Let $z,k$ be positive integers, and let $\vec{\kappa}$ be a tuple of length $z+k+1$ such that
  • $\kappa_{i}=\aleph_{0}$ for $i<z$, and
  • for $i\geq z$ the $\kappa_{i}$ are uncountable and strictly increasing.
Then $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$.
Note this shows $H^{n}(X(\aleph_{0},\aleph_{1},\ldots,\aleph_{n+1})^{-},\underline{\mathbb{Z}_{2}})\neq 0$ in ZFC.
To apply Theorem C, we let $B=\{z,z+1,\ldots,z+k\}$, so that $X(\vec{\kappa}^{B})=\alpha(\aleph_{0})^{z}$, the cartesian product of $z$ many copies of the one-point compactification of $\aleph_{0}$. We define $\varphi:X(\vec{\kappa}^{B})\to Z^{k}(\vec{\kappa}\upharpoonright_{B})$ by
[…]
(Here, by “even” we mean “of the form $\omega\cdot\lambda+2k$ for $\lambda$ limit, $k<\omega$”. This can be replaced with any sequence of unbounded, co-unbounded subsets of the $\kappa_{i}$.) Notice this is well-defined even if any subset of the $m_{i}$ are set to $\infty$; thus $\varphi$ is continuous. Moreover, if not every $m_{i}$ is $\infty$, the resulting map is mod-finite constant in the $\beta_{0}$-direction, and is thus a coboundary. So, modulo Theorem C, we need only show that the map
$\varphi(\infty,\ldots,\infty)$ $:\quad\left(\prod_{z\leq i\leq z+k}\kappa_{i}\right)\to\mathbb{Z}_{2}$
[…]
is nontrivial as a top cocycle in $\vec{\kappa}$.
To do this, we define properties $P_{j}$ for $j\leq k$ which may or may not hold of functions $f:\prod_{i\leq j}\kappa_{z+i}\to\mathbb{Z}_{2}$, inductively as follows:
  • A function $f:\kappa_{z}\to\mathbb{Z}_{2}$ has property $P_{0}$ if and only if it’s constant on a cofinite subset of $\kappa_{z}$.
  • For $j>0$, a function $f:\prod_{i\leq j}\kappa_{z+i}\to\mathbb{Z}_{2}$ has property $P_{j}$ if and only if there is $\gamma^{*}\in\kappa_{z+j}$ such that for any $\gamma_{0},\gamma_{1}\geq\gamma^{*},f(-,\gamma_{0})+f(-,\gamma_{1})$ has property $P_{j-1}$.
We first show that for $j\geq 1$, if $f:\prod_{i\leq j}\kappa_{z+i}\to\mathbb{Z}_{2}$ is a coboundary, then $f$ has property $P_{j}$:
  • If $j=1$, then $f:\kappa_{z}\times\kappa_{z+1}\to\mathbb{Z}_{2}$ is such that there are $g_{z},g_{z+1}$ such that $g_{z}+g_{z+1}=f$ and $g_{z+i}$ is mod-finite constant in the $i$-th coordinate. Then for every $\beta\in\kappa_{z},$ the set
    $\{\gamma\in\kappa_{z+1}:g_{z+1}(\beta,\gamma)\neq g_{z+1}(\beta,\infty)\}\subseteq\kappa_{z+1}$
    is finite; since $\kappa_{z}<\kappa_{z+1}$, we can find $\gamma^{*}\in\kappa_{z+1}$ above all these finite sets. Put $\gamma_{0},\gamma_{1}\geq\gamma^{*}$ and take $B\subseteq\kappa_{z}$ cofinite such that, for any $\beta\in B$ we have
    $g_{z}(\beta,\gamma_{0})=g_{z}(\infty,\gamma_{0})\text{ and }g_{z}(\beta,\gamma_{1})=\gamma_{z}(\infty,\gamma_{1}).$
    Then for any $\beta\in B$ we have
    $\phantom{=i}f(\beta,\gamma_{0})+f(\beta,\gamma_{1})$
    $=g_{z}(\beta,\gamma_{0})+g_{z+1}(\beta,\gamma_{0})+g_{z}(\beta,\gamma_{1})+g_{z+1}(\beta,\gamma_{1})$
    $=g_{z}(\infty,\gamma_{0})+g_{z+1}(\beta,\infty)+g_{z}(\infty,\gamma_{1})+g_{z+1}(\beta,\infty)$
    $=g_{z}(\infty,\gamma_{0})+g_{z}(\infty,\gamma_{1})$
    which is independent of $\beta$. So $f(-,\gamma_{0})+f(-,\gamma_{1})$ is constant on a cofinite set and thus has property $P_{1}$.
  • If $j>1$, suppose $f:\prod_{i\leq j}\kappa_{z+i}\to\mathbb{Z}_{2}$ is a coboundary, so $f=\sum_{i\leq j}g_{z+i}$ where $g_{z+i}$ is mod-finite constant in the $i$-th coordinate. In this case, for every $\vec{\beta}\in\prod_{i<j}$, the set
    $\{\gamma\in\kappa_{j}:g_{z+j}(\vec{\beta},\gamma)\neq g_{z+j}(\vec{\beta},\infty)\}$
    is again finite, so since $\kappa_{z+j}$ is strictly larger than each other $\kappa_{z+i}$ we can find $\gamma^{*}\in\kappa_{z+j}$ such that $\gamma\geq\gamma^{*}$ implies $g_{z+j}(-,\gamma)=g_{z+j}(-,\infty)$. Let $\gamma_{0},\gamma_{1}\geq\gamma^{*}$. Then we get
    $\phantom{=i}f(-,\gamma_{0})+f(-,\gamma_{1})$
    $=\left(\sum_{i<j}g_{z+i}(-,\gamma_{0})\right)+g_{z+j}(-,\gamma_{0})+\left(\sum_{i<j}g_{z+i}(-,\gamma_{1})\right)+g_{z+j}(-,\gamma_{1})$
    $=\sum_{i<j}\left(g_{z+i}(-,\gamma_{0})+g_{z+i}(-,\gamma_{1})\right),$
    which exhibits $f(0,\gamma_{0})+f(-,\gamma_{1})$ as a coboundary. By induction this means $f(0,\gamma_{0})+f(-,\gamma_{1})$ has property $P_{j-1}$ so that $f$ has property $P_{j}$.
It thus remains to show that our map $\varphi(\infty,\ldots,\infty)$ does not have property $P_{k}$. We again do this by induction on $k$:
  • If $k=1$ then $\varphi(\infty,\ldots,\infty)(\beta,\gamma)=1$ if and only if $\beta_{0}<\aleph_{0}$ and $\gamma$ is even. Fixing any $\gamma^{*}$, take $\gamma_{0},\gamma_{1}>\gamma^{*}$ even and odd, respectively. Then $\varphi(\infty,\ldots,\infty)(\beta,\gamma_{0})+\varphi(\infty,\ldots,\infty)(\beta,\gamma_{1})=1$ if and only if $\beta<\aleph_{0}$; since $\kappa_{z}>\aleph_{0}$, this is not constant on a cofinite subset of $\kappa_{z}$.
  • Now, suppose $\varphi(\infty,\ldots,\infty)$ had property $P_{j}$, with witness $\gamma^{*}\in\kappa_{z+j}$. Again choose $\gamma_{0},\gamma_{1}\geq\gamma^{*}$ even and odd respectively, and notice that
    $\varphi(\infty,\ldots,\infty)(\vec{\beta},\gamma_{0})+\varphi(\infty,\ldots,\infty)(\vec{\beta},\gamma_{1})=\varphi(\infty,\ldots,\infty)(\vec{\beta}).$
    The left hand side has property $P_{j-1}$ by hypothesis, but the left side does not, by induction. Thus $\varphi(\infty,\ldots,\infty)$ does not have property $P_{j}$.
We conclude by Theorem C that $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$.  ∎
Proposition 5.6 is typical of the intended application of Theorem C: namely, to reduce an intermediate cohomology computation to a question in top cohomology, which is more easily tractable with elementary inductions like those in the proof.
6. Further work and conclusion
Chief among the outstanding questions is whether the cardinal arithmetic assumptions of section 3 are necessary, and indeed whether any other independence phenomena arise:
Question 6.1.
Does ZFC decide the vanishing/nonvanishing of $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ for all $\vec{\kappa}$?
Additional questions about the $X(\vec{\kappa})^{-}$ are important in applications as well. Aoki shows that the cup product structure on the cohomology ring of
$X(\aleph_{0},\aleph_{1})^{-}\times\cdots\times X(\aleph_{2n,2n+1})^{-}$
admits a nonzero product of size $n+1$, and asks if there is a smaller locally profinite set with this property. Further work is required to determine if the machinery introduced here can be applied to resolve this question.
Lastly, as we have not resolved all cohomology groups of the $X(\vec{\kappa})^{-}$, it is worth pointing out some small cases that remain unsolved:
Question 6.2.
Which of the following are consistent?
  • $H^{1}(X(\aleph_{1},\aleph_{2},\aleph_{2})^{-},\underline{\mathbb{Z}_{2}})=0$? $H^{1}(X(\aleph_{1},\aleph_{2},\aleph_{2})^{-},\underline{\mathbb{Z}_{2}})\neq 0$?
  • $H^{1}(X(\aleph_{2},\aleph_{2},\aleph_{2})^{-},\underline{\mathbb{Z}_{2}})=0$? $H^{1}(X(\aleph_{2},\aleph_{2},\aleph_{2})^{-},\underline{\mathbb{Z}_{2}})\neq 0$?
  • $H^{1}(X(\aleph_{1},\aleph_{2},\aleph_{3})^{-},\underline{\mathbb{Z}_{2}})=0$? (consistently nonzero if $2^{\aleph_{1}}\geq\aleph_{3}$)
  • $H^{2}(X(\aleph_{0},\aleph_{0},\aleph_{0},\aleph_{1})^{-},\underline{\mathbb{Z}_{2}}=0$? $H^{2}(X(\aleph_{0},\aleph_{0},\aleph_{0},\aleph_{1})^{-},\underline{\mathbb{Z}_{2}}\neq 0$?